A 4.0 mol sample of O2 gas starts off with a volume of 20 L at 270 K. It is allowed to undergo an adiabatic expansion against a constant pressure of 600 torr until the volume has increased by a factor of 3. Calculate q, w, ∆T, ∆U, and ∆H, for this process. Note: consider this process carefully. Is it reversible? Is the final pressure 600 torr? Treat the gas as ideal.
Answer #1

In an adiabatte process q=0 W = - Pex DV W = - 600 Torr X (38202 - 202.) -600 TOXY (602-202) = -60oTon X 402 = - 600 Toro x 1W ST in((p-R) OT = – 3200 3 de mol (29.4 I - 8.314 ] kmol knol, AT = -37.9k LAT 2 - 38k DU SU = qtw AU = 04 (-3.2 kJ) (DU ;-3DH = DU + URAT = - 32003 t de mol 88.31% 3 -384. & mol DH ² - 4 4 619 OH 2-4500J q=0 W =-3200 J IT = -38k AU= -3200 T OH = -2